### v1

@@ 8 Conclusion
We have formalized a universal Diophantine construction that implies a stronger version of Hilbert’s Tenth Problem. This construction leads to the novel Theorem 1.1 and, moreover, it has been developed largely in parallel to its Isabelle formalization. In the course of the project, we also built a substantial library of advanced facts about multivariate polynomials and formalized results from a textbook on number theory. The total development consists of 20,000 lines of code.
@@
The formalization was executed by a student workgroup at a time when only a very early draft of the underlying mathematics was available. Hence we say that this draft manuscript was formalized in-situ, in its natural environment—a mathematics department—, and before a version of publication quality had been prepared.
Interacting with a proof assistant significantly shaped the development of the mathematical text: we refined arguments, removed or weakened redundant assumptions, and corrected various minor bugs. For example, we had overlooked a dependency on a lemma in the literature [18, Lemma 10] which only became apparent in the formalization process. We found that for certain technical parts of the proof, Isabelle functioned as a true proof assistant: supporting us authors by keeping track of assumptions and bounds so that we could focus on the more conceptual elements of the proof.
@@
Future research could extend this work by formalizing universal pairs over ${\mathbb{N}}$. We believe that the infrastructure we have built in this project would provide a strong starting point for such work, eliminating the need to formalize further prerequisites.
Since this project bridges the frontier between mathematical research and formalization in a proof assistant, further progress will naturally also involve new mathematical discoveries. Hilbert’s Tenth Problem has been proven to be solvable for $\delta\leq 2$ or alternatively $\nu=1$ [8], but remains completely open for $\delta=3$ and largely open for the whole range $2\leq\nu\leq 10$.

### v2

@@ 8 Conclusion
We have formalized a universal Diophantine construction that implies a stronger version of Hilbert’s Tenth Problem. Already the pre-print [2] of our novel Theorem 1.1 thus appears with the certificate of correctness presented herein. This result has been largely developed in parallel to its Isabelle formalization. In the course of the project, we also built a substantial library of advanced facts about multivariate polynomials and formalized results from a textbook on number theory. The total development consists of 20,000 lines of code.
@@
The formalization was executed by a student workgroup at a time when only a very early draft of the underlying mathematics was available. Hence we say that this draft manuscript was formalized in-situ, in its natural environment – a mathematics department – , and before a version of publication quality had been prepared.
Interacting with a proof assistant significantly shaped the development of the mathematics and the mathematical text: we refined arguments, removed or weakened redundant assumptions, and corrected a large number of minor bugs throughout our original draft. Even our theorems benefited significantly from the formalization. In fact, it was not before we completed its formalization that we managed to obtain the correct formula for $\eta(\nu,\delta)$, now given in Theorem˜1.1. Also the first pre-print of this article still contained a wrong formula.
To give another example, we had also overlooked that one of our technical lemmas depended on a lemma in the literature [18, Lemma 10]. Without noticing our initial fallacious proof of the technical lemma, the team responsible for its formalization had rediscovered a correct proof on their own. This subtle but severe omission in our original proof was discovered only during the final review phase of our manuscript. We then used our already complete formalization to write a correct pen-and-paper proof. Thanks to the structured proof language of Isabelle, even our collaborators who had no experience with Isabelle were able to easily read the formal proof.
Besides those two major mistakes, we found that for many technical parts of the proof, Isabelle functioned as a true proof assistant: supporting us authors by keeping track of assumptions and bounds so that we could focus on the more conceptual elements of the proof. For a complete account of the benefits and challenges of this unusual method of doing mathematics, we refer to the extended reflection in our full article [2, Section 7].
@@
Future research could extend this work by formalizing universal pairs over $\mathbb{N}$} as constructed in the article by Jones [8]. We believe that the infrastructure we have built in this project would provide a strong starting point for such work, eliminating the need to formalize further prerequisites on multivariate polynomials.
Since this project bridges the frontier between mathematical research and formalization in a proof assistant, further progress will naturally also involve new mathematical discoveries. Hilbert’s Tenth Problem has been proven to be solvable for $\delta\leq 2$ or alternatively $\nu=1$ [8], but remains largely open for $\delta=3$ and for the whole range $2\leq\nu\leq 10$.